OpenStudy (anonymous):

Equal masses of 2 liquids are mixed together. The liquids have densities of p and 2p. Assuming that there is no change in volume, determine the density of the mixture.

OpenStudy (anonymous):

let V volume of each liquid be taken.if the mass of less dense liquid be M then the mass of more dense liquid is 2M.so total of 2+1=3M is the mass of mixture and 2V is the volume of mixture. therefore the density of mixture is =3M/2V= 1.5p

OpenStudy (anonymous):

the answer shud be 4/3 p :/

OpenStudy (anonymous):

oh.. just saw that it is 'Equal masses' not equal 'vol'... let M be the mass of each liquid.. volume of lighter one= M/p vol of denser liq = M/2p therefore net density= total mass/total vol = 2M/(M/p + M/2p) = 4/3 p